OpenStudy (angela210793):

Find this one.... {[1+(x-1)/2]/1-1/x}-x/(x-1) --------------------------- x/2 I really hope u'll get it

OpenStudy (anonymous):

Thank you for the star.

OpenStudy (angela210793):

Uw :)...any idea abt this 1 ?

OpenStudy (anonymous):

Let me give it a try

OpenStudy (angela210793):

Thnx ^_^ :)

OpenStudy (dumbcow):

\[\frac{x-2}{x-1}\]

OpenStudy (angela210793):

@dumbcow Will you please explain me how did u get that? Please...

OpenStudy (dumbcow):

just do it step by step, combine fractions and use property of dividing fractions (Flip and multiply)

OpenStudy (anonymous):

\[\frac{((1+(x-1)/2)/1-1/x)-x/(x-1)}{\frac{x}{2}}=\frac{((-2+x) (-1+x))}{ x^2} \]

OpenStudy (angela210793):

thnx :)

OpenStudy (anonymous):

In a rush to get out a solution I did not verify the result. ie: x=17 shows that the left side and the right side are not equal. Recalculated and it appears that at first glance that the fraction is equal to one.

OpenStudy (angela210793):

it's ok :) I'll check it tomorrow :):):) Thnx anyway

OpenStudy (toxicsugar22):

dumbcow can u help me

OpenStudy (anonymous):

The fraction is equal to 1. The problem fraction with the braces and brackets changed to parentheses:\[\frac{((1 + (x - 1)/2)/1 - 1/x) - x/(x - 1)}{\frac{x}{2}} \]Multiply the numerator by the reciprocal of the denominator adding surrounding parentheses to the numerator.\[\frac{2}{x}(((1 + (x - 1)/2)/1 - 1/x) - x/(x - 1)) \]The following is a step by step sequence of fraction modifications, mainly simplifications to individual terms.\[\frac{2}{x}\left(((1+(x-1)/2)/1-1/x)-\frac{x}{-1+x}\right) \]\[\frac{2}{x}\left(\left(\left.\frac{1+x}{2}\right/1-1/x\right)-\frac{x}{-1+x}\right) \]\[\frac{2}{x}\left(\left(\frac{1+x}{2}/\frac{-1+x}{x}\right)-\frac{x}{-1+x}\right) \]\[\frac{2}{x}\left(\frac{x (1+x)}{2 (-1+x)}-\frac{x}{-1+x}\right) \]Multiply through by 2/x\[\left(\frac{2}{x}\right)\left(\frac{x (1+x)}{2 (-1+x)}\right)-\left(\frac{x}{-1+x}\right)\left(\frac{2}{x}\right) \]Expand the products on each side of the minus sign and simplify.\[\frac{1}{-1+x}+\frac{x}{-1+x}-\frac{2}{-1+x} \]\[\frac{1+x-2}{-1+x} \]\[\frac{-1+x}{-1+x}=1 \]

OpenStudy (angela210793):

Wooooow...Thanks a loooooooooooooooooooooooooooot!!!!!!!!! :) ^_^ Thanks :)